The ratio by mass of sulfur dioxide to oxygen in a chemical reactor is 5:3. If the total mass of both compounds is 72 grams, what is the mass of oxygen.
Answer by stanbon(75887)

The ratio by mass of sulfur dioxide to oxygen in a chemical reactor is 5:3. If the total mass of both compounds is 72 grams, what is the mass of oxygen.
-----------------------
5:3 is the same as 5x to 3x where oxygen = 3x
-----
Equation:
5x + 3x = 72 g
8x = 72
x = 9
----
Ans: oxygen = 3x = 3*9 = 27 grams
-----------
Cheers,
Stan H.
------------